A carpet cleaning machine holds 40litres of water. It is filled with water at 15 ∘ C and the water is heated by a 3.0kW element to a temperature of 70 ∘C.
How much energy is added to the water?
Mass of 1.0litre=1.0kg; s.h.c.=4200J/(kgK)]
1
Expert's answer
2020-09-21T08:32:09-0400

The enegy transfered to the water is given by the following expression:


"Q = cm(T_2 - T_1)"

where "c = 4200J/(kg\cdot K)" is the s.h.c. of water, "m = 40kg" is the mass of the water, "T_1 = 15\degree C = 288K" is the initial temperature of the water and "T_2 = 70\degree C = 343K" is the final temperature of the water. Finally, obtain:


"Q = 4200\cdot 40\cdot (343-288) = 9240000J = 9.24\times10^6J = 9.24MJ"

Answer. 9.24*10^6 J or 9.24 MJ.